# Transmon Qubit Constraints on Dark Matter-Nucleon Scattering

### v1

@@ Abstract
Direct detection experiments are currently the main detection effort for dark matter scattering on Earth. As the dark matter wind impinges on Earth’s surface, dark matter particles can deposit energy inside direct detection experiments, which search for anomalous recoils of their Standard Model (SM) targets. This provides a robust lab-based test of the dark matter-SM scattering cross section. However, direct detection experiments are limited to dark matter masses above about a GeV for standard analyses assuming nuclear recoils [1], or MeV-scale masses when exploiting the Migdal effect [2, 3, 4, 5, 6] or electron recoils [7, 8, 9].
To date, no definitive dark matter induced recoil signal has been detected, motivating a push for new approaches to search for dark matter scattering outside the known sensitivity of direct detection experiments [10]. To probe lighter dark matter, or dark matter with less energy that that coming from the dark matter wind, lower threshold devices are required. New proposals have been made in recent years on multiple fronts, with new detection possibilities with superconductors [11, 12, 13, 14, 15, 16], superfluids [17, 18, 19], polar crystals [20, 21, 22], topological materials [23], and Dirac materials [24, 25, 26, 27, 28].
@@ I.3 Testing Dark Matter Sub-Components
Sub-component dark matter can be motivated from the perspective that the dark sector may be vast, with more than one particle candidate, not dissimilar to the spectrum of particles observed within the SM. There are a range of particle dark matter models considered in the literature that naturally can provide dark matter candidates in our parameter space of interest, including dark matter candidates with quantum chromodynamics (QCD) charge such as the sexaquark [48, 49, 50, 51, 43]. Another example explored in $e.g.$ Ref. [37] is the scenario where asymmetric dark matter is coupled to the dark photon, with predictions in previously untested parameter space to which we are sensitive. In general, sub-fractions of the full dark matter density may be expected when the interaction rate is fairly large, leading to large annihilation rates in the early Universe and therefore a depletion in the relic dark matter abundance.
@@ II.2 Calculation of Scattering Rates and Energy Deposition
Here, as we have an aluminium target, $\bar{\omega}_{d}=25.5$ meV. As an example benchmark, for thermalized dark matter with 1 GeV mass and a maximum momentum transfer of $q=20$ keV, the typical phonon number is therefore $n=0.29$.
@@ II.3 Dark Matter Induced Quasiparticle Density
where $\bar{\Gamma}=40\,\mathrm{s^{-1}\mu m^{3}}$ for aluminium. The quasiparticle generation rate $\bar{\Gamma}$ has some associated uncertainty, and so when making plots we will take $\overline{\Gamma}=4$ or $400\,\mathrm{s^{-1}}\,\mu\mathrm{m}^{3}$ for conservative and optimistic values respectively.
@@ III New Dark Matter Parameter Space Sensitivities
Figure 3 shows the bounds and sensitivities we derive for spin-independent scattering using a transmon qubit, using the incoming Galactic dark matter, under the assumption the dark matter makes up the full relic abundance ($i.e.$ $f_{\chi}=1$). This is in direct one to one correspondence with the usual direct detection bounds, that is, all lines here have about the same assumptions of the local dark matter density and velocity. The limits in Fig. 3 can be simply rescaled linearly in $f_{\chi}$ for subcomponent dark matter. We see that our bounds on halo dark matter are stronger than lab-based experiments below about 100 MeV, as we already had shown in Ref. [29]. We also show new projected sensitivities for a few benchmark scenarios, assuming improved future single quasiparticle devices. We have taken a quasiparticle density about an order of magnitude than our existing transmon qubit, as labeled as the dashed line with $n_{\rm qp}=10^{-3}\,\mu\mathrm{m}^{-3}$, at two quasiparticle temperatures of 180 mK and 50 mK as labeled. As noted in the previous section, these choices are motivated as the 180 mK line is effectively the setup of Ref. [59], except recombination limited, and 50 mK is chosen as another benchmark as it corresponds to approaching the limit of what can be realistically achieved in the near future. Note that all estimates assume recombination-limited quasiparticle density modeling. Active work to understand sources of quasiparticle poisoning for qubit error mitigation, both from qubit drive [63] and radiation [64, 65], will contribute to even more stringent sensitivities than the projections shown, as quasiparticle backgrounds are reduced further.
@@
Figure 4: New constraints on the dark matter-nucleon spin-independent scattering cross section $\sigma_{\chi N}^{\rm SI}$ as a function of dark matter mass $m_{\chi}$, using transmon qubit measurements for the Earth-bound dark matter population, for varying fractions of the dark matter density $f_{\chi}$ as labeled. Dashed line is the conservative bound, solid is the optimistic bound (see text for details). Complementary limits are also shown as labeled.
Figure 4 shows our new limits on the dark matter-nucleon spin-independent scattering cross section as a function of dark matter mass, using transmon qubit measurements for the Earth-bound dark matter population, for varying fractions of the dark matter density $f_{\chi}$ as labeled. The solid line corresponds to the optimistic bound based on quasiparticle generation rate uncertainties, and the dashed line is conservative. We see that while the $f_{\chi}=1$ case (dark matter candidate corresponds to the full dark matter density) overlaps most with existing experiments, as soon as this density is decreased other experiments weaken or disappear, while our bounds are not largely changed in comparison.
In Fig. 4, note that we linearly rescale the complementary direct detection limits only from below; the ceilings of the limits also decrease but non-linearly with decreasing dark matter fraction, and require a dedicated calculation [77]. The reason the ceilings scale non-linearly with decreasing dark matter fraction, is that the dark matter velocity is also relevant when the scattering limit is near the detection threshold [77]. As only CRESST has had its limit ceilings recast for subfractions [77], we adopt the recast CRESST bounds for both the lower and upper ends of the bounds, but only rescale the other direct detection constraints from below. From the point of view of how much new parameter space our transmon qubit probes, this is a conservative choice, although the ceiling limit change with decreasing dark matter subfractions in any case is not very large [77]. In the case where $f_{\chi}=0.1$, already the astrophysical bounds disappear. This is because limits relying on structure formation require the dark matter candidate to make up most of the dark matter content, and as soon as dark matter is a smaller component, it has a negligible effect on structure.
@@ IV Summary and Outlook
For the Galactic halo dark matter population, transmon qubits provide constraints on dark matter-nucleon scattering stronger than direct detection for dark matter masses below about 100 MeV. We showed the future promise for transmon qubits or any low-quasiparticle density device to detect light dark matter, considering both lower quasiparticle densities, as well as devices cooled to lower temperatures. We expect that low quasiparticle density devices can reach cross sections as low as about $10^{-31}$ cm2 for dark matter masses below about 100 MeV for Galactic halo dark matter, and many orders of magnitude below this for thermalized dark matter. Active work to understand sources of quasiparticle poisoning for qubit error mitigation, both from qubit drive [63] and radiation [64, 65] will led to reduced quasiparticle backgrounds, and even better sensitivities in the future as quasiparticle backgrounds are reduced further. Overall, new low-quasiparticle density devices as they are actively created for applications such as quantum computing, infrared telescopes, or radiation detectors, offer substantial potential to probe dark matter with low energy deposition, and are a promising path forward to detect light dark matter.

### v2

@@ Abstract
Direct detection experiments are currently the main detection effort for dark matter scattering on Earth. As the dark matter wind impinges on Earth’s surface, dark matter particles can deposit energy inside direct detection experiments, which search for anomalous recoils of their Standard Model (SM) targets. This provides a robust lab-based test of the dark matter-SM scattering cross section. However, direct detection experiments are limited to dark matter masses above about a GeV for analyses assuming nuclear recoils [1], or MeV-scale masses when exploiting the Migdal effect [2, 3, 4, 5, 6, 7] or electron recoils [8, 9, 10, 11].
To date, no definitive dark matter induced recoil signal has been detected, motivating a push for new approaches to search for dark matter scattering outside the known sensitivity of direct detection experiments [12]. To probe lighter dark matter, or dark matter with less energy than that coming from the dark matter wind, lower threshold devices are required. New proposals have been made in recent years on multiple fronts, with new detection possibilities with superconductors [13, 14, 15, 16, 17, 18, 19, 20], superfluids [21, 22, 23], polar crystals [24, 25, 26], doped semiconductors [27], topological materials [28], Dirac materials [29, 30, 31, 32, 33], boosted dark matter in various detectors [34, 35, 36, 37, 38, 39, 40, 41], and crystal defects [42].
@@ I.3 Testing Dark Matter Sub-Components
Sub-component dark matter can be motivated from the perspective that the dark sector may be vast, with more than one particle candidate – after all, the visible sector contains a large number of particles within the SM. There are a range of particle dark matter models considered in the literature that naturally can provide dark matter candidates in our parameter space of interest, including dark matter candidates with quantum chromodynamics (QCD) charge such as the sexaquark [62, 63, 64, 65, 57]. Another example explored in $e.g.$ Ref. [51] is the scenario where asymmetric dark matter is coupled to the dark photon, with predictions in previously untested parameter space to which our setup has sensitivity. In general, sub-fractions of the full dark matter density may be expected when the interaction rate is fairly large, leading to large annihilation rates in the early Universe and therefore a depletion in the relic dark matter abundance.
@@ II.2 Calculation of Scattering Rates and Energy Deposition
Here, as we have an aluminium target, $\bar{\omega}_{d}=25.5$ meV and $m_{d}=27\,\mathrm{GeV}$. As an example benchmark, for thermalized dark matter with 1 GeV mass and a maximum momentum transfer of $q=20$ keV, the typical phonon number is therefore $n=0.29$.
@@ II.3 Dark Matter Induced Quasiparticle Density
where $\bar{\Gamma}=40\,\mathrm{s^{-1}\mu m^{3}}$ for aluminium [71]. The quasiparticle generation rate $\bar{\Gamma}$ has some associated uncertainty due to the limitations of the quasiparticle diffusion model, and so when making plots we will take $\overline{\Gamma}=4$ or $400\,\mathrm{s^{-1}}\,\mu\mathrm{m}^{3}$ for conservative and optimistic values respectively.
@@ III New Dark Matter Parameter Space Sensitivities
Figure 3 shows the bounds and sensitivities we derive for spin-independent scattering using a transmon qubit, using the incoming Galactic dark matter, under the assumption the dark matter makes up the full relic abundance ($i.e.$ $f_{\chi}=1$). This is in direct one to one correspondence with the usual direct detection bounds, that is, all lines here have about the same assumptions of the local dark matter density and velocity. The limits in Fig. 3 can be simply rescaled linearly in $f_{\chi}$ for subcomponent dark matter. We see that our bounds on halo dark matter are stronger than lab-based experiments below about 100 MeV, as we already had shown in Ref. [43]. We also show new projected sensitivities for a few benchmark scenarios, assuming improved future single quasiparticle devices. We have taken a quasiparticle density about an order of magnitude smaller than our existing transmon qubit, as labeled as the dashed line with $n_{\rm qp}=10^{-3}\,\mu\mathrm{m}^{-3}$, at two quasiparticle temperatures of 180 mK and 50 mK as labeled. As noted in the previous section, these choices are motivated as the 180 mK line is effectively the setup of Ref. [73], except recombination limited, and 50 mK is chosen as another benchmark as it corresponds to approaching the limit of what can be realistically achieved in the near future. Note that all estimates assume recombination-limited quasiparticle density modeling. Active work to understand sources of quasiparticle poisoning for qubit error mitigation, both from qubit drive [77] and radiation [78, 79], will contribute to even more stringent sensitivities than the projections shown, as quasiparticle backgrounds are reduced further.
@@
Figure 4: New constraints on the dark matter-nucleon spin-independent scattering cross section $\sigma_{\chi N}^{\rm SI}$ as a function of dark matter mass $m_{\chi}$, using transmon qubit measurements for the Earth-bound dark matter population, for varying fractions of the dark matter density $f_{\chi}$ as labeled. Solid line is the conservative bound, dashed is the optimistic bound (see text for details). Complementary limits are also shown as labeled.
Figure 4 shows our new limits on the dark matter-nucleon spin-independent scattering cross section as a function of dark matter mass, using transmon qubit measurements for the Earth-bound dark matter population, for varying fractions of the dark matter density $f_{\chi}$ as labeled. The dashed line corresponds to the optimistic bound based on quasiparticle generation rate uncertainties, and the solid line is conservative. We see that while the $f_{\chi}=1$ case (dark matter candidate corresponds to the full dark matter density) overlaps most with existing experiments, as soon as this density is decreased other experiments weaken or disappear, while our bounds are not largely changed in comparison.
In Fig. 4, note that we linearly rescale the complementary direct detection limits only from below; the ceilings of the limits also decrease but non-linearly with decreasing dark matter fraction, and require a dedicated calculation [91]. The reason the ceilings scale non-linearly with decreasing dark matter fraction, is that the dark matter velocity is also relevant when the scattering limit is near the detection threshold [91]. As only CRESST has had its limit ceilings recast for subfractions [91], we adopt the recast CRESST bounds for both the lower and upper ends of the bounds, but only rescale the other direct detection constraints from below. From the point of view of how much new parameter space our transmon qubit probes, this is a conservative choice, although the ceiling limit change with decreasing dark matter subfractions in any case is not very large [91]. In the case where $f_{\chi}=0.1$, the astrophysical bounds weaken or disappear. This is because limits relying on structure formation require the dark matter candidate to make up most of the dark matter content, and as soon as dark matter is a smaller component, it has a negligible effect on structure.
@@ IV Summary and Outlook
For the Galactic halo dark matter population, transmon qubits provide constraints on dark matter-nucleon scattering stronger than direct detection for dark matter masses below about 100 MeV. We showed the future promise for transmon qubits or any low-quasiparticle density device to detect light dark matter, considering both lower quasiparticle densities, as well as devices cooled to lower temperatures. We expect that low quasiparticle density devices can reach cross sections as low as about $10^{-31}$ cm2 for dark matter masses below about 100 MeV for Galactic halo dark matter, and many orders of magnitude below this for thermalized dark matter. Active work to understand sources of quasiparticle poisoning for qubit error mitigation, both from qubit drive [77] and radiation [78, 79] will lead to reduced quasiparticle backgrounds, and even better sensitivities in the future as quasiparticle backgrounds are reduced further. Overall, new low-quasiparticle density devices as they are actively created for applications such as quantum computing, infrared telescopes, or radiation detectors, offer substantial potential to probe dark matter with low energy deposition, and are a promising path forward to detect light dark matter.